# Are Graph Neural Networks Optimal Approximation Algorithms?

### v5

@@
Abstract
In this work we design graph neural network architectures that capture optimal approximation algorithms for a large class of combinatorial optimization problems, using powerful algorithmic tools from semidefinite programming (SDP). Concretely, we prove that polynomial-sized message-passing algorithms can represent the most powerful polynomial time algorithms for Max Constraint Satisfaction Problems assuming the Unique Games Conjecture. We leverage this result to construct efficient graph neural network architectures, OptGNN, that obtain high-quality approximate solutions on landmark combinatorial optimization problems such as Max-Cut, Min-Vertex-Cover, and Max-3-SAT. Our approach achieves strong empirical results across a wide range of real-world and synthetic datasets against solvers and neural baselines. Finally, we take advantage of OptGNN’s ability to capture convex relaxations to design an algorithm for producing bounds on the optimal solution from the learned embeddings of OptGNN.
@@ 1 Introduction
Combinatorial Optimization (CO) is the class of problems that optimize functions subject to constraints over discrete search spaces. They are often NP-hard to solve and to approximate, owing to their typically exponential search spaces over nonconvex domains. Nevertheless, their important applications in science and engineering (Gardiner et al., 2000; Zaki et al., 1997; Smith et al., 2004; Du et al., 2017) has engendered a long history of study rooted in the following simple insight. In practice, CO instances are endowed with domain-specific structure that can be exploited by specialized algorithms (Hespe et al., 2020; Walteros & Buchanan, 2019; Ganesh & Vardi, 2020). In this context, neural networks are natural candidates for learning and then exploiting patterns in the data distribution over CO instances.
The emerging field at the intersection of machine learning (ML) and combinatorial optimization (CO) has led to novel algorithms with promising empirical results for several CO problems. However, similar to classical approaches to CO, ML pipelines have to manage a tradeoff between efficiency and optimality. Indeed, prominent works in this line of research forego optimality and focus on parametrizing heuristics (Li et al., 2018; Khalil et al., 2017; Yolcu & Póczos, 2019; Chen & Tian, 2019) or by employing specialized models (Zhang et al., 2023; Nazari et al., 2018; Toenshoff et al., 2019; Xu et al., 2021; Min et al., 2022) and task-specific loss functions (Amizadeh et al., 2018; Karalias & Loukas, 2020; Wang et al., 2022; Karalias et al., 2022; Sun et al., 2022). Exact ML solvers that can guarantee optimality often leverage general techniques like branch and bound (Gasse et al., 2019; Paulus et al., 2022) and constraint programming (Parjadis et al., 2021; Cappart et al., 2019), which offer the additional benefit of providing approximate solutions together with a bound on the distance to the optimal solution. The downside of those methods is their exponential worst-case time complexity. Striking the right balance between efficiency and optimality is quite challenging, which leads us to the central question of this paper:
@@
To answer this question, we build on the extensive literature on approximation algorithms and semidefinite programming. Convex relaxations of CO problems via semidefinite programming are the fundamental building block for breakthrough results in the design of efficient algorithms for NP-hard combinatorial problems, such as the Goemans-Williamson approximation algorithm for Max-Cut (Goemans & Williamson, 1995) and the use of the Lovász theta function to find the maximum independent set on perfect graphs (Lovász, 1979; Grötschel et al., 1981). For several problems, it is known that if the Unique Games Conjecture (UGC) is true, then the approximation guarantees obtained through semidefinite programming relaxations are indeed the best that can be achieved (Raghavendra, 2008; Barak & Steurer, 2014). We will leverage these results to provide an affirmative answer to our question. Our contributions can be organized into theory and experiments.
The theoretical contribution of this work is to show that a polynomial time message-passing algorithm approximates the solution of an SDP with the optimal integrality gap for the class of Maximum Constraint Satisfaction Problems (Max-CSP), assuming the UGC. Max-CSP is the problem of given a set of constraints over variables satisfy as many as possible. The key theoretical insight is that a message-passing algorithm can be used to compute gradient updates for the penalized Lagrangian of an overparameterized reformulation of the SDP in (Raghavendra, 2008). This in turn leads to our main contribution, OptGNN, a GNN architecture that generalizes our message-passing algorithm and therefore captures its approximation guarantee. Leveraging this result, we provide a perturbation analysis demonstrating OptGNN is efficiently PAC-learnable.
Our second contribution is empirical. We show that our theoretical construction can be used directly to design efficient graph neural network pipelines for CO that are easy to implement and train. By training with the SDP objective as a loss function, OptGNN learns embeddings that can be regarded as feasible fractional solutions of an overparameterized low-rank SDP, which are subsequently rounded to feasible integral solutions. We show OptGNN achieves strong empirical results against solvers and neural baselines across a broad battery of datasets for landmark CO problems such as Max-Cut, Vertex-Cover, and Max-3-SAT.
@@
Figure 1: Schematic representation of OptGNN. During training, OptGNN produces node embeddings $\mathbf{v}$ using message passing updates on the graph $G$. These embeddings are used to compute the penalized objective $\mathcal{L}_{p}(\mathbf{v};G)$. OptGNN is trained by minimizing the average loss over the training set. At inference time, the fractional solutions (embeddings) $\mathbf{v}$ for an input graph $G$ produced by OptGNN are rounded using randomized rounding.
Finally, to underscore the fact that OptGNN captures powerful convex relaxations, we construct dual certificates of optimality, i.e bounds that are provably correct, from OptGNN embeddings for the Max-Cut problem that are virtually tight for small synthetic instances. See our discussion on general neural certification schemes for extracting dual certificates from OptGNN networks in Appendix C.2. To summarize, the contribution of this paper is twofold:
@@
  • We construct a polynomial time message passing algorithm for solving the SDP of Raghavendra (2008) for the broad class of maximum constraint satisfaction problems (including Max-Cut, Max-SAT, etc.), that is optimal barring the possibility of significant breakthroughs in the foundations of algorithms.
@@
  • We construct a graph neural architecture to capture this message-passing algorithm and show that it achieves strong results against solvers and neural baselines.
@@ 2 Background and Related Work
Optimal Approximation Algorithms. It is typically difficult to prove that an algorithm achieves the best approximation guarantee for a given problem, as it is hard to rule out the existence of a more powerful algorithm. The Unique Games Conjecture (UGC) (Khot, 2002) is a striking development in the theory of approximation algorithms because it can circumvent precisely this obstacle. If true, it implies approximation hardness results which often match the approximation guarantees of the best-known algorithms (Raghavendra & Steurer, 2009b; Raghavendra et al., 2012). For that reason, those algorithms are also sometimes referred to as UGC-optimal. More importantly, the UGC implies something even stronger: there is a general algorithm based on semidefinite programming that achieves the best possible approximation guarantees for the broad class of Max Constraint Satisfaction Problems (Raghavendra, 2008). For a complete exposition on the topic of UGC and approximation algorithms, we refer the reader to Barak & Steurer (2014). Our theoretical contribution builds on these ideas to construct a neural architecture that is a candidate optimal approximation algorithm.
@@
Neural approximation algorithms and their limitations. The design of neural approximation algorithms typically involves the use of state of the art architectures like graph neural networks (GNN). There has been important progress in characterizing the combinatorial capabilities of those architectures, including bounds on the approximation guarantees achievable by GNNs on bounded degree graphs (Sato et al., 2019; 2021) and conditional impossibility results for solving classic combinatorial problems such as Max-Independent-Set and Min-Spanning-Tree (Loukas, 2019). It has been shown that a GNN can implement a distributed local algorithm that straightforwardly obtains a 1/2-approximation for Max-SAT (Liu et al., 2021), which is also achievable through a simple randomized algorithm (Johnson, 1973). Recent work proves there are barriers to the approximation power of GNNs for combinatorial problems including Max-Cut and Min-Vertex-Cover (Gamarnik, 2023) for constant depth GNNs. Other approaches to obtaining approximation guarantees propose avoiding the dependence of the model on the size of the instance with a divide-and-conquer strategy. This is done by partitioning the input into smaller subproblems of bounded size and using the neural network to solve each subproblem (McCarty et al., 2021; Kahng et al., 2023).
Convex Relaxations and Machine Learning. Convex relaxations are the workhorse of approximation algorithms; a discrete problem is first relaxed, solved, and then rounded to produce a high quality solution. SDPs are a particularly powerful tool in this context, as they often yield the strongest approximation guarantees for combinatorial problems.
In this work we show how SDP-based approximation algorithms can be incorporated into the architecture of neural networks. Efficient methods have been developed for solving low rank SDPs Burer & Monteiro (2003); Wang et al. (2017); Wang & Kolter (2019); Boumal et al. (2020). Building on these ideas, we design and then embed a UGC-optimal message passing algorithm within a GNN architecture. Beyond approximation algorithms, there is work on designing differentiable Max-SAT solvers via SDPs to facilitate symbolic reasoning in neural architectures (Wang et al., 2019). This approach uses a fixed algorithm for solving a novel SDP relaxation for Max-SAT, and aims to learn the structure of the SAT instance. In our case, the instance is given, but our algorithm is learnable, and we seek to predict the solution.
Semidefinite programming has a myriad of applications in machine learning including the design of high-dimensional differentiable extensions of discrete functions (Karalias et al., 2022) for neural CO, the use of neural networks to solve feasibility SDPs in quantum information tasks Kriváchy et al. (2021), global minimization of functions via a kernel approach to semidefinite programming (Rudi et al., 2020), and leveraging the connection between SVMs and the Lovász theta function to efficiently find the largest common dense subgraph among a collection of graphs (Jethava et al., 2013).
@@
Convex relaxations are also instrumental in Mixed Integer Programming. Convex relaxation guided branch-and-bound tree search is guaranteed to terminate with optimal integral solutions to Mixed Integer Linear Programs (MILP). Proposals for incorporating neural networks into the MILP pipeline include providing a “warm start” (Benidis et al., 2023) to the solver, learning branching heuristics (Gasse et al., 2019; Nair et al., 2020; Gupta et al., 2020; Paulus et al., 2022), and learning cutting plane protocols Paulus et al. (2022). A recent line of work studies the capabilities of neural networks to solve linear programs (Chen et al., 2022; Qian et al., 2023). It is shown that GNN’s can represent LP solvers, which may in turn explain the success of learning branching heuristics (Qian et al., 2023). In a similar line of work, neural nets are used to learn branching heuristics for CDCL SAT solvers (Selsam & Bjørner, 2019; Kurin et al., 2020; Wang et al., 2021). Finally, convex optimization has also found applications Numeroso et al. (2023) in the neural algorithmic reasoning paradigm (Veličković et al., ; Veličković et al., 2022) where neural networks are trained to solve problems by learning to emulate discrete algorithms in higher dimensional spaces.
@@
Learning frameworks for CO. There are numerous viable design paradigms when it comes to developing effective neural CO architectures. Some works employ a supervised approach and rely on learning from execution traces of expert algorithms or labeled solutions to the problem they are trying to solve (Li et al., 2018; Selsam et al., 2018; Prates et al., 2019; Vinyals et al., 2015; Joshi et al., 2019; 2020; Gasse et al., 2019; Ibarz et al., 2022; Georgiev et al., 2023). Obtaining labels for combinatorial problems can be computationally costly. Consequently, a broad class of algorithms for neural CO involves training without any access to labels or partial solutions. This includes approaches based on Reinforcement Learning (Ahn et al., 2020; Böther et al., 2022; Barrett et al., 2020; 2022; Bello et al., 2016; Khalil et al., 2017; Yolcu & Póczos, 2019; Chen & Tian, 2019), and other self-supervised methods (Brusca et al., 2023; Karalias et al., 2022; Karalias & Loukas, 2020; Tönshoff et al., 2022; Schuetz et al., 2022a; b; Amizadeh et al., 2019; Dai et al., 2020; Sun et al., 2022; Wang et al., 2022; Amizadeh et al., 2018; Gaile et al., 2022). For a complete overview of the field, we refer the reader to the relevant survey papers (Cappart et al., 2023; Bengio et al., 2021). Our work falls into the latter category since only the problem instance is sufficient for training and supervision signals are not required.
@@ 3 Optimal Approximation Algorithms with Neural Networks
We begin by showing that solving the Max-Cut problem using a vector (low-rank SDP) relaxation and a simple projected gradient descent scheme amounts to executing a message-passing algorithm on a graph. This sets the stage for our main result for the class of maximum constraint satisfaction problems (Max-CSP), which generalizes this insight. We reformulate the UGC-optimal SDP for Max-CSP in (SDP 1 top of pg. 19). Our main Theorem 3.1 exhibits a message passing algorithm (Algorithm 1) for solving SDP 1. We then capture Algorithm 1 via a message passing architecture (OptGNN) with learnable weights. Thus, by construction OptGNN captures algorithms with UGC-optimal approximation guarantees for Max-CSP. Furthermore, we prove that OptGNN is efficiently PAC-learnable (see Lemma 3.1) as a step towards explaining its empirical performance.
@@ 3.1 Solving Combinatorial Optimization Problems with Message Passing
$\max_{(x_{1},x_{2},...,x_{N})}\quad$ $\sum_{(i,j)\in E}\frac{1}{2}(1-x_{i}x_{j})$
subject to: $x_{i}^{2}=1\quad\forall i\in[N]$
@@
The global optimum of the integer program is the Max-Cut. Noting that discrete variables are not amenable to the tools of continuous optimization, a standard technique is to ’lift’ the quadratic integer problem and solve it with a rank constrained SDP. Here we introduce a matrix $X\in\mathbb{R}^{N\times N}$ of variables, where we index the $i$’th row and $j$’th column entry as $X_{ij}$.
$\max_{X}\quad$ $\sum_{(i,j)\in E}\frac{1}{2}(1-X_{ij})$ (1)
subject to: $X_{ii}=1\quad\forall i\in[N]$
$X\succeq 0$
$\text{rank}(X)=r$
The intuition is that a rank $r=1$ solution to Problem 1 is equivalent to solving the integer program. A common approach is to replace the integer variables $x_{i}$ with vectors $v_{i}\in\mathbb{R}^{r}$ and constrain $v_{i}$ to lie on the unit sphere.
@@
$\min_{v_{1},v_{2},\dots,v_{N}}\quad$ $-\sum_{(i,j)\in E}\frac{1}{2}(1-\langle v_{i},v_{j}\rangle)$ (2)
subject to: $\|v_{i}\|=1\quad\forall i\in[N]$
$v_{i}\in\mathbb{R}^{r}$ (3)
@@
This problem can be solved with a fast iterative algorithm Burer & Monteiro (2003). The landscape of this nonconvex optimization can be benign in that all local minima are approximately global minima (Ge et al., 2016) and variations on stochastic gradient descent converge to its optimum (Bhojanapalli et al., 2018; Jin et al., 2017) under a variety of smoothness and compactness assumptions. Specifically, for large enough $r$ Boumal et al. (2020); O’Carroll et al. (2022), simple algorithms such as block coordinate descent (Erdogdu et al., 2019) can find an approximate global optimum of the objective. Projected gradient descent is a natural approach for solving the minimization problem in equation 2. In iteration $t$ (for $T$ iterations), update vector $v_{i}$ as
$\hat{v}_{i}^{t+1}$ $=v_{i}^{t}-\eta\sum\nolimits_{j\in N(i)}v_{j}^{t}$ (4)
$v_{i}^{t+1}$ $={\hat{v}_{i}^{t+1}}/{\|\hat{v}_{i}^{t+1}\|},$ (5)
@@
where $\eta\in\mathbb{R}^{+}$ is an adjustable step size and $N(i)$ the neighborhood of node $i$. The gradient updates to the vectors are local, i.e., each vector is updated by aggregating information from its neighboring vectors. Hence, projected gradient descent can be implemented as a message-passing algorithm (See our precise definition of message passing algorithm in Definition Definition). Subsequently, the vectors can be rounded to an integral solution by taking a random projection with a hyperplane or any of a variety of rounding methods.
@@
OptGNN for Max-Cut. Our main contribution in this paper builds on the following observation. We may generalize the dynamics described above by considering an overparametrized version of the gradient descent updates in equations 4 and 5. Let $M_{1},M_{2},...,M_{T}\in\mathbb{R}^{r\times 2r}$ be a set of $T$ learnable matrices corresponding to $T$ layers of a neural network. Then for layer $t$ and embedding $v_{i}$ we have
$\hat{v}_{i}^{t+1}$ :$\vcentcolon=M_{t}\left(\begin{bmatrix}v_{i}^{t}\\ \sum_{j\in N(i)}v_{j}\end{bmatrix}\right)$ (8)
$v_{i}^{t+1}$ :$\vcentcolon={\hat{v}_{i}^{t+1}}/{\|\hat{v}_{i}^{t+1}\|},$ (9)
More generally, we can write the dynamics as
$\hat{v}_{i}^{t+1}$ :$\vcentcolon=M_{t}\left(\text{AGG}(v_{i}^{t},\{v_{j}^{t}\}_{j\in N(i)})\right)$ (10)
$v_{i}^{t+1}$ :$\vcentcolon=\text{NONLINEAR}(\hat{v}_{i}^{t+1}),$ (11)
where AGG is a function of the node embedding and its neighboring embeddings. It is straightforward to see that these update dynamics can be captured by a Graph Neural Network (GNN). For a Max-Cut instance $\Lambda$ specified by a graph $G_{\Lambda}$, we refer to equation 8 and equation 9 as the OptGNN for Max-Cut with learnable parameters $M=\{M_{t}\}_{t\in[T]}$ over a constraint graph $G_{\Lambda}$ denoted $\text{OptGNN}_{(M,\Lambda)}$ (for formal definition see Definition Definition). Furthermore, we exhibit a message passing algorithm 2 that generalizes the dynamics of 4 and 5 to the entire class of Max-CSPs, which provably converges in polynomial iterations for a reformulation of the canonical SDP relaxation of Raghavendra (2008) (see SDP 1 top of pg. 19 for reformulated SDP). For our main theorem on convergence see the informal statement Theorem 3.1 and formal statement Theorem C.1. For ease of reading, we suggest the reader first inspect the informal description of algorithm 1 before the formal description in algorithm 2. We also extend our OptGNN construction to the Min-Vertex-Cover problem and Max-3-SAT (see derivation in Appendix A and Appendix B).

### v7

@@
Are Graph Neural Networks Optimal Approximation Algorithms?
Morris Yau
MIT CSAIL
morrisy@mit.edu &Eric Lu
Harvard
ericlu01@g.harvard.edu &Nikolaos Karalias
MIT CSAIL
stalence@mit.edu &Jessica Xu
Reverie Labs
jessica.peng.xu@gmail.com &Stefanie Jegelka
MIT CSAIL
stefje@mit.edu
Equal Contribution.
@@ 1 Introduction
Combinatorial Optimization (CO) is the class of problems that optimize functions subject to constraints over discrete search spaces. They are often NP-hard to solve and to approximate, owing to their typically exponential search spaces over nonconvex domains. Nevertheless, their important applications in science and engineering (gardiner2000graph; zaki1997new; smith2004design; du2017multi) has engendered a long history of study rooted in the following simple insight. In practice, CO instances are endowed with domain-specific structure that can be exploited by specialized algorithms (hespe2020wegotyoucovered; walteros2019maximum; ganesh2020unreasonable). In this context, neural networks are natural candidates for learning and then exploiting patterns in the data distribution over CO instances.
The emerging field at the intersection of machine learning (ML) and combinatorial optimization (CO) has led to novel algorithms with promising empirical results for several CO problems. However, similar to classical approaches to CO, ML pipelines have to manage a tradeoff between efficiency and optimality. Indeed, prominent works in this line of research forego optimality and focus on parametrizing heuristics (li2018combinatorial; khalil2017learning; yolcu2019learning; chen2019learning) or by employing specialized models (zhang2023let; nazari2018reinforcement; toenshoff2019run; xu2021reinforcement; min2022can) and task-specific loss functions (amizadeh2018learning; karalias2020erdos; wang2022unsupervised; karalias2022neural; sun2022annealed). Exact ML solvers that can guarantee optimality often leverage general techniques like branch and bound (gasse2019exact; paulus2022learning) and constraint programming (parjadis2021improving; cappart2019improving), which offer the additional benefit of providing approximate solutions together with a bound on the distance to the optimal solution. The downside of those methods is their exponential worst-case time complexity. Striking the right balance between efficiency and optimality is quite challenging, which leads us to the central question of this paper:
@@
{wrapfigure}
@@
L0.6 [Uncaptioned image] Schematic representation of OptGNN. During training, OptGNN produces node embeddings $\mathbf{v}$ using message passing updates on the graph $G$. These embeddings are used to compute the penalized objective $\mathcal{L}_{p}(\mathbf{v};G)$. OptGNN is trained by minimizing the average loss over the training set. At inference time, the fractional solutions (embeddings) $\mathbf{v}$ for an input graph $G$ produced by OptGNN are rounded using randomized rounding.
To answer this question, we build on the extensive literature on approximation algorithms and semidefinite programming (SDP) which has led to breakthrough results for NP-hard combinatorial problems, such as the Goemans-Williamson approximation algorithm for Max-Cut (goemans1995improved) and the use of the Lovász theta function to find the maximum independent set on perfect graphs (lovasz1979shannon; grotschel1981ellipsoid). For several problems, it is known that if the Unique Games Conjecture (UGC) is true, then the approximation guarantees obtained through SDP relaxations are indeed the best that can be achieved (raghavendra2008optimal; barak2014sum). The key insight of our work is that a polynomial time message-passing algorithm (see Definition Definition) approximates the solution of an SDP with the optimal integrality gap for the class of Maximum Constraint Satisfaction Problems (Max-CSP), assuming the UGC. This algorithm can be naturally parameterized to build a graph neural network, which we call OptGNN.
Our contributions can be summarized as follows:
@@
  • We construct a polynomial-time message passing algorithm (see Definition Definition) for solving the SDP of raghavendra2008optimal for the broad class of maximum constraint satisfaction problems (including Max-Cut, Max-SAT, etc.), that is optimal barring the possibility of significant breakthroughs in the field of approximation algorithms.
@@
  • We construct a graph neural network architecture which we call OptGNN, to capture this message-passing algorithm. We show that OptGNN is PAC-learnable with polynomial sample complexity.
  • We describe how to generate optimality certificates from the learned representations of OptGNN i.e., provable bounds on the optimal solution.
  • Empirically, OptGNN is simple to implement (see pseudocode in Appendix 5) and we show that it achieves competitive results on 3 landmark CO problems and several datasets against classical heuristics, solvers and state-of-the-art neural baselines  [^1] Code is available at: https://github.com/penlu/bespoke-gnn4do.. Furthermore, we provide out-of-distribution (OOD) evaluations and ablation studies for OptGNN that further validate our parametrized message-passing approach.
@@ 2 Background and Related Work
Optimal Approximation Algorithms. Proving that an algorithm achieves the best approximation guarantee for NP-hard problems is an enormous scientific challenge as it requires ruling out that better algorithms exist (i.e., a hardness result). The Unique Games Conjecture (UGC) (khot2002power) is a striking development in the theory of approximation algorithms because it addresses this obstacle. If true, it implies approximation hardness results for a large number of hard combinatorial problems that match the approximation guarantees of the best-known algorithms (raghavendra2009towards; raghavendra2012reductions). For that reason, those algorithms are also sometimes referred to as UGC-optimal. More importantly, the UGC implies that for all Max-CSPs there is a general UGC-optimal algorithm based on semidefinite programming (raghavendra2008optimal). For a complete exposition on the topic of UGC and approximation algorithms, we refer the reader to barak2014sum.
@@
Neural approximation algorithms and their limitations. There has been important progress in characterizing the combinatorial capabilities of modern deep learning architectures, including bounds on the approximation guarantees achievable by GNNs on bounded degree graphs (sato2019approximation; sato2021random) and conditional impossibility results for solving classic combinatorial problems such as Max-Independent-Set and Min-Spanning-Tree (loukas2019graph). It has been shown that a GNN can implement a distributed local algorithm that straightforwardly obtains a 1/2-approximation for Max-SAT (liu2021can), which is also achievable through a simple randomized algorithm (johnson1973approximation). Recent work proves there are barriers to the approximation power of GNNs for combinatorial problems including Max-Cut and Min-Vertex-Cover (gamarnik2023barriers) for constant depth GNNs. Other approaches to obtaining approximation guarantees propose avoiding the dependence of the model on the size of the instance with a divide-and-conquer strategy; the problem is subdivided into smaller problems which are then solved with a neural network (mccarty2021nn; kahng2023nn).
@@
Convex Relaxations and Machine Learning. Convex relaxations are crucial in the design of approximation algorithms. In this work, we show how SDP-based approximation algorithms can be incorporated into the architecture of neural networks. We draw inspiration from the algorithms that are used for solving low-rank SDPs (burer03; wang2017imposing; wang2019low; boumal2020deterministic). Beyond approximation algorithms, there is work on designing differentiable Max-SAT solvers via SDPs to facilitate symbolic reasoning in neural architectures (wang2019satnet). This approach uses a fixed algorithm for solving a novel SDP relaxation for Max-SAT, and aims to learn the structure of the SAT instance. In our case, the instance is given, but our algorithm is learnable, and we seek to predict the solution. SDPs have found numerous applications in machine learning including neural network verification (brown2022unified), differentiable learning with discrete functions (karalias2022neural), kernel methods (rudi2020finding; jethava2013lovasz) and quantum information tasks (krivachy2021high). Convex relaxation are also instrumental in integer programming where branch-and-bound tree search is guaranteed to terminate with optimal integral solutions to Mixed Integer Linear Programs (MILP). Proposals for incorporating neural networks into the MILP pipeline include providing a “warm start” (benidis2023solving) to the solver, learning branching heuristics (gasse2019exact; nair2020solving; gupta2020hybrid; paulus2022learning), and learning cutting plane protocols (paulus2022learning). A recent line of work studies the capabilities of neural networks to solve linear programs (chen2022representing; qian2023exploring). It is shown that GNNs can represent LP solvers, which may in turn explain the success of learning branching heuristics (qian2023exploring). In a similar line of work, neural nets are used to learn branching heuristics for CDCL SAT solvers (selsam2019guiding; kurin2020can; wang2021neurocomb). Finally, convex optimization has also found applications (numeroso2023dual) in the neural algorithmic reasoning paradigm (velivckovicneural; velivckovic2022clrs) where neural networks are trained to solve problems by learning to emulate discrete algorithms in higher dimensional spaces.
@@
Learning frameworks for CO. A common approach to neural CO is to use supervision either in the form of execution traces of expert algorithms or labeled solutions (li2018combinatorial; selsam2018learning; prates2019learning; vinyals2015pointer; joshi2019efficient; joshi2020learning; gasse2019exact; ibarz2022generalist; georgiev2023neural). Obtaining labels for combinatorial problems can be computationally costly which has led to the development of neural network pipelines that can be trained without access to labels or partial solutions. This includes approaches based on Reinforcement Learning (ahn2020learning; bother2022s; barrett2020exploratory; barrett2022learning; bello2016neural; khalil2017learning; yolcu2019learning; chen2019learning), and other self-supervised methods (brusca2023maximum; karalias2022neural; karalias2020erdos; tonshoff2022one; schuetz2022combinatorial; schuetz2022graph; amizadeh2019pdp; dai2020framework; sun2022annealed; wang2022unsupervised; amizadeh2018learning; gaile2022unsupervised). Our work falls into the latter category since only the problem instance is sufficient for training and supervision signals are not required. For a complete overview of the field, we refer the reader to the relevant survey papers (cappart2023combinatorial; bengio2021machine).
@@ 3 Optimal Approximation Algorithms with Neural Networks
We begin by showing that solving the Max-Cut problem using a vector (low-rank SDP) relaxation and a simple projected gradient descent scheme amounts to executing a message-passing algorithm on the input graph. We then generalize this insight to the entire class of Max-CSPs. We reformulate the UGC-optimal SDP for Max-CSP in SDP 1. Our main Theorem 3.1 exhibits a message passing algorithm (Algorithm 1) for solving SDP 1. We then capture Algorithm 1 via a message passing GNN with learnable weights (see Definition Definition for definition of Message Passing GNN). Thus, by construction OptGNN captures algorithms with UGC-optimal approximation guarantees for Max-CSP. Furthermore, we prove that OptGNN is efficiently PAC-learnable (see Lemma 3.1) as a step towards explaining its empirical performance.
@@ 3.1 Solving Combinatorial Optimization Problems with Message Passing
$\max_{x_{1},x_{2},...,x_{N}}\quad$ $\sum_{(i,j)\in E}\frac{1}{2}(1-x_{i}x_{j})\quad\text{subject to:}\quad x_{i}^{2}=1\quad\forall i\in[N]$
@@
The global optimum of the integer program is the Max-Cut. Noting that discrete variables are not amenable to the tools of continuous optimization, a standard technique is to ’lift’ the quadratic integer problem: replace the integer variables $x_{i}$ with vectors $v_{i}\in\mathbb{R}^{r}$ and constrain $v_{i}$ to lie on the unit sphere.
@@
$\min_{v_{1},v_{2},\dots,v_{N}}\quad$ $-\sum_{(i,j)\in E}\frac{1}{2}(1-\langle v_{i},v_{j}\rangle)\quad\text{subject to:}\quad\|v_{i}\|=1\quad\forall i\in[N]\quad v_{i}\in\mathbb{R}^{r}.$ (1)
@@
This nonconvex relaxation of the problem admits an efficient algorithm burer03. Furthermore, all local minima are approximately global minima (GLMspurious16) and variations of stochastic gradient descent converge to its optimum (bhojanapalli2018smoothed; jin2017escape) under a variety of smoothness and compactness assumptions. Specifically, for large enough $r$ (boumal2020deterministic; o2022burer), simple algorithms such as block coordinate descent (erdogdu2019convergence) can find an approximate global optimum of the objective. Projected gradient descent is a natural approach for solving the minimization problem in equation 1. In iteration $t$ (for $T$ iterations), update vector $v_{i}$ as
${v}_{i}^{t+1}$ $=\text{NORMALIZE}\left(v_{i}^{t}-\eta\sum\nolimits_{j\in N(i)}v_{j}^{t}\right),$ (2)
@@
where NORMALIZE enforces unit Euclidean norm, $\eta\in\mathbb{R}^{+}$ is an adjustable step size, and $N(i)$ the neighborhood of node $i$. The gradient updates to the vectors are local, each vector is updated by aggregating information from its neighboring vectors (i.e., it is a message-passing algorithm).
@@
OptGNN for Max-Cut. Our main contribution in this paper builds on the following observation. We may generalize the dynamics described above by considering an overparametrized version of the gradient descent updates in equations 2. Let $M_{1},M_{2},...,M_{T}\in\mathbb{R}^{r\times 2r}$ be a set of $T$ learnable matrices corresponding to $T$ layers of a neural network. Then for layer $t$ and embedding $v_{i}$ we define a GNN update
${v}_{i}^{t+1}$ :$\vcentcolon=\text{NORMALIZE}\left(M_{t}\left(\begin{bmatrix}v_{i}^{t}\\ \sum_{j\in N(i)}v_{j}^{t}\end{bmatrix}\right)\right)$ (3)
More generally, we can write the dynamics as ${v}_{i}^{t+1}\vcentcolon=\text{NORMALIZE}(M_{t}(\text{AGG}(v_{i}^{t},\{v_{j}^{t}\}_{j\in N(i)}))$, where AGG is a function of the node embedding and its neighboring embeddings. We will present a message passing algorithm 1 that generalizes the dynamics of 2 to the entire class of Max-CSPs (see example derivations in Appendix A and Appendix B), which provably converges in polynomial iterations for a reformulation of the canonical SDP relaxation of raghavendra2008optimal (see SDP 1). Parametrizing this general message-passing algorithm will lead to the definition of OptGNN (see Definition Definition).